, Laird realizes he hates mathematics and his Job, and decides to jump off of a building. If his height off the ground at time t is given by h(t)=-2t^2+30t+500 , at what time will he reach his maximum height, and what will it be?
Answer by Alan3354(69443)

h(t)=-2t^2+30t+500
Max height will be at the vertex of the parabola.
t = -b/2a = -30/-4 = 7.5 seconds
---------------------------------
h(7.5) = -2*56.25 + 30*7.5 + 500
max h = 612.5 feet
-------------
It seems odd to jump off a building and ascend 112 feet. Maybe he was trying to jump over it.